OpenStudy (cutiecomittee123):

Write an equation of the line that is parallel to 2x-3y=5 and passes through the point (2,-1)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What do parallel lines have in common?

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

Do you know how to read the slope off of the equation of a line?

OpenStudy (cutiecomittee123):

I figured it out. It is 2x-3y=7

OpenStudy (skullpatrol):

$$\Huge 2x-3y=5$$ $$\Huge\text{Want the slope-intercept form:}$$ $$\Huge{y = mx + b}$$